### 06 Apr 2026 v2

# An angular-momentum preserving dissipative model for the point-mass $N$-body problem

Abstract
A simple mathematical model emulating energy dissipation due to tidal effects is proposed. In this model, forces acting between masses remove energy but preserve the total angular momentum of the system. We study the effect of such forces on the particular family of orbits in central configurations, and show that a specific dependence on the mutual distances between the bodies leads to homographic equations equivalent to those of the two-body problem with dissipation. We then describe in detail the topology of solutions of the dissipative two-body system via Poincaré compactification. Finally, we present equations averaged over Keplerian motion showing no influence of the dissipation on periapsis precession.
keywords:
Celestial Mechanics, $N$-body problem, Dissipative systems
1 Introduction
When predicting the fate of $N$-bodies mutually attracted by gravity, two main frameworks are often used to model their dynamical evolution. In the conservative one, orbits are either regular (periodic or quasi-periodic) or chaotic, where bodies may escape, collide, or revolve around each other indefinitely [Smale, Contopoulos_book, Ferraz-Mello_book]. In the dissipative one, these scenarios can still occur, but pairs of bounded bodies, given sufficient time, eventually end up in spin-orbit coupling, facing each other while moving in circular orbits, or collide [Darwin].
In fact, the spin-orbit coupling among bounded pairs extends to systems with any number of bodies ($N>2$), as the whole system is progessively reduced, pair by pair, to one of the final states described: escapes, collisions, or synchronizations [Scheeres_2012].
In detail, one can start by assuming the conservative dynamics for $N$-bodies, generally taking place on the hypersurface $E:\mathbb{R}^{3N}\times T\mathbb{R}^{3N}\to\mathbb{R}$, corresponding to the system’s mechanical energy
$E(\textbf{r},\dot{\textbf{r}})=T(\dot{\textbf{r}})+U(\textbf{r}),$ (1)
for $\textbf{r}=\left(\textbf{r}_{1}\;...\;\textbf{r}_{N}\right)=\left(x_{1}\;y_{1}\;z_{1}\;...\;x_{N}\;y_{N}\;z_{N}\right)\in\mathbb{R}^{3N}$ as the cartesian position for all $N$-bodies, and $\dot{\textbf{r}}=\frac{d\textbf{r}}{dt}\in T\mathbb{R}^{3N}$ as their velocities. The kinetic energy
$T(\dot{\textbf{r}})=\frac{1}{2}\sum_{i=1}^{N}m_{i}\left(\dot{\textbf{r}}_{i}\cdot\dot{\textbf{r}}_{i}\right)$ (2)
and the gravitational potential energy
$U(\textbf{r})=-G\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\frac{m_{i}m_{j}}{|\textbf{r}_{ij}|},$ (3)
are computed for masses $m_{i}$ with relative distances $\textbf{r}_{ij}=\textbf{r}_{i}-\textbf{r}_{j}\in\mathbb{R}^{3}$.
If no external forces act on the system, and regardless of internal constraints, the dynamics will preserve both the energy function (1) and the total angular momentum
$\textbf{L}=\sum_{i=1}^{N}m_{i}\left(\textbf{r}_{i}\times\dot{\textbf{r}}_{i}\right),$ (4)
due to the central nature of the forces from the gravitational potential (3). The expressions (1) to (4) hold for a reference frame fixed at and moving with the center of mass of the whole system, satisfying
$\sum_{i=1}^{N}m_{i}\textbf{r}_{i}=\sum_{i=1}^{N}m_{i}\dot{\textbf{r}}_{i}=0.$ (5)
In this context, Smale [Smale] proved that for any gravitational system of point-masses, the critical points of the energy function (1) restricted to a constant angular momentum hypersurface (4) correspond to relative equilibrium configurations, i.e.,
$\frac{d|\textbf{r}_{ij}|}{dt}=0\qquad\forall i,j.$ (6)
On top of that, Moeckel [Moeckel_1990] showed that for $N>2$ any relative equilibrium is a saddle-point of the energy function, when restricted to a constant angular momentum level. Later, Moeckel [Moeckel] proved similar results for systems of interacting rigid bodies.
In the presence of tides, which dissipate energy but not angular momentum, a distinct conclusion can be derived from the results of Smale and Moeckel. For $N\geq 2$, energy dissipation leads to two possible physical scenarios: either bodies collide and/or move apart and ultimately become isolated masses; or, after successive collisions and escapes, pairs of masses remain in circular motion, where no further energy can be dissipated (see [Scheeres_2002, Scheeres_2012] for details). The combination of Smale’s and Moeckel’s results assert that, despite the existence of relative equilibria, their instability prevents systems with more than two bodies to remain in such states indefinitely.
It is well known that the behavior observed in nature is dissipative, as different mechanisms act on any group of celestial bodies due to internal friction (tidal effects), several types of drag (motion within gases, dust, plasma, etc.), and radiative effects [Ferraz-Mello_book]. However, in empty space, these dissipative forces are generally weak, and the timescales required to reach final states are often unattainable. As a result, one typically observes systems in transient configurations lasting for millions or even billions of years, closely resembling stability. This sugests that the nature of dissipative effects allows for prolonged transients near semi-equilibrium states, making them relevant to the dynamical evolution as such states likely represent most of a system’s lifetime and determine their path toward any final state it may reach.
Several models for dissipative effects have been studied [Gabriel, Wang, Celletti_2007, Celletti_2011, Margheri, Ferraz-Mello], mainly involving Poynting-Robertson drag, solar wind, atmospheric drag, or surface contact. In these cases, dissipation typically acts as a drag force opposing the body’s velocity through a damping term $\alpha\dot{\textbf{r}}$ in the equation of motion
$m\ddot{\textbf{r}}=-\alpha\dot{\textbf{r}}-\nabla U(\textbf{r}).$ (7)
Even though these models correspond to relevant and realistic scenarios, they remove both energy and angular momentum from the system. By contrast, at planetary scale, the most prominent dissipative effect is gravitational in origin, arising from tides, which preserves angular momentum [Goldreich, Meyer].
In this context, we introduce a simple model for dissipation between celestial bodies that, while dissipating energy, preserves angular momentum. This is achieved by introducing a central damping force that breaks time symmetry while maintaining rotational symmetry.
We remark that a force similar, though not identical, to the one we propose has already appeared in a PhD thesis by Chong Ai under the supervision of Mark Levi [Chong_Ai]. Since their force has the same physical features as the one we propose here, their qualitative results for the dissipative Kepler problem are similar to ours. They also studied dissipative effects in the restricted three-body problem, an investigation that we are also pursuing. To our knowledge, the exact force expression we propose, which under dissipation leaves invariant a manifold that in the conservative case is generated by the homographic solutions, has not been proposed before.
In what follows, we present the model and its main properties in Section 2. Section 3 discusses its effects on planar central configurations, while Section 4 and 5 provide topological and numerical results for the simpler case of two bodies – results that also apply to planar central motions for any number of bodies.
2 Model
To incorporate energy dissipation while conserving angular momentum, we draw on Mignard’s tidal force approximation for two nearly spherical bodies (eq. (5) in [Mignard]). We simplify his expression to a point-mass scenario by neglecting rotational terms, yielding the dissipative force $\textbf{f}^{\,ij}_{\textrm{diss}}$ between a pair of masses $m_{i}$ and $m_{j}$
$\textbf{f}^{\,ij}_{\textrm{diss}}(\dot{\textbf{r}}_{ij},\textbf{r}_{ij})=-\alpha\frac{Gm_{i}m_{j}}{|\textbf{r}_{ij}|^{d}}\left(\dot{\textbf{r}}_{ij}\cdot\hat{\textbf{r}}_{ij}\right)\hat{\textbf{r}}_{ij},$ (8)
where $\hat{\textbf{r}}_{ij}=\frac{\textbf{r}_{i}-\textbf{r}_{j}}{|\textbf{r}_{i}-\textbf{r}_{j}|}\in\mathbb{R}^{3}$ is the versor connecting the $i$-th and $j$-th bodies. The dissipation strength $\alpha$ is constant and has units of $\left[L^{d-3}T\right]$. Physically, it reflects rheological and viscoelastic effects [^1] In principle, the dissipation constant $\alpha$ may vary from body to body, so that $\textbf{f}^{\,ij}_{\textrm{diss}}(\dot{\textbf{r}}_{ij},\textbf{r}_{ij})=-\alpha_{i}\alpha_{j}\frac{Gm_{i}m_{j}}{|\textbf{r}_{ij}|^{d}}\left(\dot{\textbf{r}}_{ij}\cdot\hat{\textbf{r}}_{ij}\right)\hat{\textbf{r}}_{ij},$ In a forthcoming paper, focusing on the three-body problem, we use this formulation. Here we assume a uniform $\alpha$ for all bodies. This distinction is irrelevant in the case of two bodies. . Since the current study focuses on a mathematical approach, we treat the radial dependence exponent $d$ as a free parameter, as it will be discussed further in the text. A direct comparison to Mignard’s model would yield $d=8$.
The force (8) is pairwise anti-symmetric ($\textbf{f}_{ij}=-\textbf{f}_{ji}$), dissipative because it depends on the relative velocity $(\dot{\textbf{r}}_{ij}=\dot{\textbf{r}}_{i}-\dot{\textbf{r}}_{j})$, and central ($\textbf{f}_{ij}\,\|\,\hat{\textbf{r}}_{ij}$). It is also homogeneous in the positions and velocities $\textbf{f}^{\,ij}_{\textrm{diss}}(\nu\dot{\textbf{r}}_{ij},\mu\textbf{r}_{ij})=\nu\mu^{-d}\textbf{f}^{\,ij}_{\textrm{diss}}(\dot{\textbf{r}}_{ij},\textbf{r}_{ij})$.
We start by considering the conservative Lagrangian for $N$ point masses
$\mathcal{L}=\frac{1}{2}\sum_{i=1}^{N}m_{i}\left(\dot{\textbf{r}}_{i}\cdot\dot{\textbf{r}}_{i}\right)+G\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\frac{m_{i}m_{j}}{|\textbf{r}_{ij}|},$ (9)
and introduce dissipation via the Rayleigh function
$\mathcal{D}=\frac{\alpha}{2}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\frac{Gm_{i}m_{j}}{|\textbf{r}_{ij}|^{d}}\left(\dot{\textbf{r}}_{ij}\cdot\hat{\textbf{r}}_{ij}\right)^{2},$ (10)
which modify the Euler-Lagrange equations as
$\frac{d}{dt}\left(\nabla_{\dot{\textbf{r}}_{i}}\mathcal{L}\right)=\nabla_{\textbf{r}_{i}}\mathcal{L}-\nabla_{\dot{\textbf{r}}_{i}}\mathcal{D};$ (11)
yielding the equations of motion
$m_{i}\ddot{\textbf{r}}_{i}=-\sum_{j\neq i}^{N}\frac{Gm_{i}m_{j}}{|\textbf{r}_{ij}|^{3}}\,\textbf{r}_{ij}+\sum_{j\neq i}^{N}\textbf{f}^{\,ij}_{\textrm{diss}}(\dot{\textbf{r}}_{ij},\textbf{r}_{ij}),$ (12)
for $\textbf{r}_{i}\in\mathbb{R}^{3}$ as the position of each mass $m_{i}$ relative to the center of mass.
2.1 Energy dissipation
To assess how the dissipative force (8) affects the system, we examine two conservative invariants of motion, namely the energy and angular momentum. Because the force is central, it immediately follows that the angular momentum (4) will remain conserved.
As for the energy, we can evaluate its dissipation by taking the scalar product of the equation of motion (12) with the velocity $\dot{\textbf{r}}$
$\begin{split}M\ddot{\textbf{r}}\cdot\dot{\textbf{r}}+\nabla_{\textbf{r}}U(\textbf{r})\cdot\dot{\textbf{r}}&=-2\mathcal{D}(\dot{\textbf{r}},\textbf{r})\end{split}$ (13)
and identifying the left hand side as the derivative of the total energy $E=T+U$, and therefore
$\frac{dE}{dt}=-2\mathcal{D}\leq 0.$ (14)
Dissipation will only cease when $\mathcal{D}=0$, which exclusively occurs in configurations of relative equilibria, or rotational central configurations, where the whole system rotates with all pairs $\textbf{r}_{ij}$ fixed.
3 Central configurations
In this section we introduce some necessary concepts on central configurations (CCs) and analyze the effect of the dissipative force (8) on this family of orbits when restricted to the planar case. For further read on CCs, we direct the reader to Moeckel’s work (Chap. 2 in [Moeckel_book]) on which we base our calculations.
Given a set of $N$-bodies with masses $m_{1},\dots,m_{N}$, a CC is an arrangement where the force on each body points towards the center of mass of the system, that is
$\nabla U(\textbf{r})+\lambda\textbf{M}\left(\textbf{r}-\textbf{c}\right)=0,$ (15)
for some constant $\lambda\in\mathbb{R}$. Here $\textbf{r}\in\mathbb{R}^{3N}$ (or $\mathbb{R}^{2N}$) is the global state vector with all position coordinates for the $N$-bodies. The mass matrix
$\textbf{M}=\text{diag}_{N}\left(\textbf{M}_{i}\right)=\begin{pmatrix}\mathbf{M}_{1}&\mathbf{0}&\cdots&\mathbf{0}\\ \mathbf{0}&\mathbf{M}_{2}&\cdots&\mathbf{0}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\cdots&\mathbf{M}_{N}\end{pmatrix}\in\mathbb{M}_{3N}\;(\text{or}\;\mathbb{M}_{2N}),$
stores the masses for each body, with $\textbf{M}_{i}=m_{i}\mathbf{I}_{3}$ (or $m_{i}\mathbf{I}_{2}$), for i=1,,Ni=1,...,N, and where $\mathbf{I}_{3}$ ($\mathbf{I}_{2}$) is the identity matrix in $\mathbb{R}^{3}$ ($\mathbb{R}^{2}$). The vector $\textbf{c}\in\mathbb{R}^{3N}$ is a stacked vector of the center of mass $\textbf{c}_{0}$, such that
$\left(\textbf{r}-\textbf{c}\right)=\left(\textbf{r}_{1}-\textbf{c}_{0}\;\;...\;\;\textbf{r}_{n}-\textbf{c}_{0}\right),$
with
$\textbf{c}_{0}=\frac{1}{M}\sum_{i=1}^{N}m_{i}\textbf{r}_{i}$
as the center of mass in 3D (or 2D); $M=\sum_{i=1}^{N}m_{i}$ is the total mass. Normally $\textbf{c}_{0}$ can be made zero if the coordinate reference frame is centered at it.
It can be shown that $\lambda=\frac{U(\textbf{r})}{I(\textbf{r})}$, where $I(\textbf{r})$ is the moment of inertia of the system
$I(\textbf{r})=\left(\textbf{r}-\textbf{c}\right)^{T}\textbf{M}\left(\textbf{r}-\textbf{c}\right)=\sum_{i=1}^{N}m_{i}\left|\textbf{r}_{i}-\textbf{c}_{0}\right|^{2}=\frac{1}{M}\sum_{i=1}^{N}\sum_{j<i}^{N}m_{i}m_{j}r_{ij}^{2}.$ (16)
3.1 Planar case
When restricted to a plane, the motion of any CC can be written in homographic form, combining a scaling of the configuration (homothetic motion) and a rotation (rigid body motion)
$\textbf{r}(t)-\textbf{c}(t)=s(t)\textbf{Q}(\theta(t))\left(\textbf{r}_{0}-\textbf{c}_{0}\right)$ (17)
where $s(t)\in\mathbb{R}$ is the scaling factor, and $\textbf{Q}(\theta(t))$ is a rotation matrix about the axis perpendicular to the plane, rotating each body uniformly by $\theta(t)$
$\textbf{Q}(\theta(t))=\textrm{diag}_{N}\left(\textbf{q}(\theta(t))\right)\in\mathbb{M}_{2N}$ (18)
where
$\textbf{q}(\theta(t))=\begin{pmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{pmatrix}\in\textrm{SO}(2);$ (19)
$(\textbf{r}_{0}-\textbf{c}_{0})\in\mathbb{R}^{2N}$ is the initial state of the CC. Note that $\textbf{c}_{0}\in\mathbb{R}^{2N}$ here generates an ambiguity with the vector $\textbf{c}_{0}\in\mathbb{R}^{2}$, denoting the position of the center of mass in the plane of motion. Although related, the ambiguity should not bring any conflict in the following calculations.
Over a homographic configuration ($\textbf{r}_{h},\dot{\textbf{r}}_{h}$) satisfying (17) the Rayleigh dissipation function (10) can be evaluated as
[…]
where $\mu_{ij}=\alpha Gm_{i}m_{j}$ and $D(r_{ij,0})=|\textbf{r}_{ij,0}|^{-d}$. When replacing the homographic form (17) into (10), the following identities were used
$\begin{split}&\frac{d\left(\textbf{r}-\textbf{c}\right)}{dt}=\left(\dot{s}(t)\textbf{Q}(\theta(t))+s(t)\textbf{Q}^{\prime}\dot{\theta}(t)\right)\left(\textbf{r}_{0}-\textbf{c}_{0}\right)\\ &\frac{d^{2}\left(\textbf{r}-\textbf{c}\right)}{dt^{2}}=\left(\ddot{s}(t)\textbf{Q}(\theta(t))+2\dot{s}(t)\textbf{Q}^{\prime}\dot{\theta}(t)+s(t)\textbf{Q}^{\prime\prime}\dot{\theta}^{2}(t)+s(t)\textbf{Q}^{\prime}\ddot{\theta}(t)\right)\left(\textbf{r}_{0}-\textbf{c}_{0}\right)\end{split}$ (21)
where
$\textbf{Q}^{\prime}=\frac{d\textbf{Q}(\theta)}{d\theta}\;\to\;\frac{d\textbf{q}}{d\theta}=\textbf{j}\textbf{q}\qquad\text{with}\qquad\textbf{j}=\textbf{q}(\pi/2),$
and
$\textbf{Q}^{\prime\prime}=\frac{d^{2}\textbf{Q}(\theta)}{d\theta^{2}}\to\frac{d^{2}\textbf{q}}{d\theta^{2}}=-\textbf{q}.$
It was also used that rotations preserve inner products and that, when restricted to the plane, they commute, implying
$\left(\textbf{j}\textbf{q}\textbf{r}_{ij,0}\right)\cdot\left(\textbf{q}\hat{\textbf{r}}_{ij,0}\right)=\left(\textbf{j}\textbf{r}_{ij,0}\right)\cdot\hat{\textbf{r}}_{ij,0}=0.$
3.2 The special case $d=3$
For $d=3$, the dissipative force $\nabla_{\dot{\textbf{r}}_{i}}\mathcal{D}$ in (20) can be expressed in terms of the gradient of the gravitational potential
$\begin{split}&\nabla_{\dot{\textbf{r}}_{i}}\mathcal{D}(\textbf{r}_{h},\dot{\textbf{r}}_{h})=\alpha\dot{s}s^{-3}\textbf{q}\nabla_{\textbf{r}_{i}}U(\textbf{r}_{0}).\end{split}$
In this case, $\alpha$ has dimensions of inverse of time. This relation between $\nabla_{\dot{\textbf{r}}_{i}}\mathcal{D}$ and $\nabla_{\textbf{r}_{i}}U$ allows us to obtain the equation for the homographic parameters ($s(t),\theta(t)$) in the presence of dissipation.
Substituting the homographic form $\textbf{r}_{h}(t)=\textbf{c}(t)+s(t)\textbf{Q}(\theta(t))(\textbf{r}_{0}-\textbf{c}_{0})$ (Eq. (17)) into the stacked equations of motion (12), and using that $\ddot{\textbf{c}}=0$, with identities (21) and multiplying by $\textbf{Q}^{-1}$, gives
[…]
where $\textbf{J}=\textrm{diag}_{N}(\textbf{j})$. To satisfy the CC condition, it must hold that
$\nabla_{\textbf{r}_{0}}U(\textbf{r}_{0})+\lambda\textbf{M}\left(\textbf{r}_{0}-\textbf{c}_{0}\right)=0,$
noticing that $\lambda=\frac{U(\textbf{r})}{I(\textbf{r})}<0$. Simultaneously, since J is a right-angle rotation, it renders $\textbf{J}(\textbf{r}_{0}-\textbf{c}_{0})$ orthogonal to $(\textbf{r}_{0}-\textbf{c}_{0})$. This splits equation (22) into terms parallel and orthogonal to $(\textbf{r}_{0}-\textbf{c}_{0})$, yielding a system of equations for the coefficients
$\begin{cases}\ddot{s}-s\dot{\theta}^{2}=\lambda\left(s^{-2}+\alpha\dot{s}s^{-3}\right)\\ s\ddot{\theta}+2\dot{s}\dot{\theta}=0\end{cases}$ (23)
Equation (23) is a dissipative Kepler-like system in the homographic parameters, with $s(t)$ playing the role of the radial coordinate, and $\theta(t)$ the angular position. Due to this equivalence, the next section treats the two-body problem under dissipation with $d=3$, providing solutions applicable both to the two-body case and to generic CCs for arbitrary $N$.
4 Two bodies
We describe the two-body system in polar coordinates $(r,\theta)$, where $v=\dot{r}$ denotes the radial velocity. We choose the mass unit such that the reduced mass is one, $m_{1}m_{2}/(m_{1}+m_{2})=1$. The equations of motion take the first-order form
$\begin{cases}\dot{r}=v\\ \dot{v}=-\partial_{r}\mathcal{H}-\partial_{v}\mathcal{D}=\frac{C^{2}}{r^{3}}-\gamma\left(\frac{1}{r^{2}}+\alpha\frac{v}{r^{3}}\right)\end{cases}$ (24)
with $\gamma=G(m_{1}+m_{2})$. The energy $\mathcal{H}$ and dissipation $\mathcal{D}$ functions are given by
$\mathcal{H}=\frac{v^{2}}{2}+\frac{C^{2}}{2r^{2}}-\frac{\gamma}{r}\qquad\text{and}\qquad\mathcal{D}=\frac{\alpha}{2}\frac{\gamma}{r^{3}}v^{2},$ (25)
and the (conserved) normalized angular momentum CC is defined by
$C=r^{2}\dot{\theta}.$ (26)
It is convenient to define the non-dimensional parameter
$c=\frac{C}{\gamma^{2/3}\alpha^{1/3}}.$ (27)
For $\alpha>0$ and in convenient time and length scales, the dynamics depends only on $c$, see equation (42).
To understand the solutions of equation (24), we proceed to a Poincaré compactification of the phase space in a series of transformations, following Dumortier et al. [Dumortier_book] (Chapter 5.1). The idea relies on mapping the infinite phase space $(r,v)\in[0,\infty)\times\mathbb{R}$ into a sphere, and analyze the dynamics at infinity via stereographic charts (planar maps) covering the sphere. This allows us to understand the asymptotic behavior of solutions as they approach infinity in the original coordinates.
4.1 Poincaré compactification
We start by multiplying the equations in (24) by $r^{3}$ to produce a polynomial vector field, which preserves the phase space portrait of the system as it is equivalent to a time reparametrization:
$\begin{cases}\dot{r}=r^{3}v\\ \dot{v}=C^{2}-\gamma\left(r+\alpha v\right),\end{cases}$ (28)
The vector field (28) has two equilibria, namely the circular orbit
$(r_{c},v_{c})=\left(\frac{C^{2}}{\gamma},0\right),$ (29)
and a second equilibrium
$(r_{e},v_{e})=\left(0,\frac{C^{2}}{\gamma\alpha}\right),$ (30)
which emerges as an artifact of the $r^{3}$ scaling and that does not exist in the conservative case for $\alpha=0$ (Fig. 1 (A) and (B)).
The eigenvalues of the vector field (28) at the circular orbit $(r_{c},v_{c})$ are
$\begin{split}a^{c}_{\pm}=-\frac{\alpha\gamma}{2}\left(1\pm\sqrt{1-\frac{4C^{6}}{\gamma^{4}\alpha^{2}}}\right)=-\frac{\alpha\gamma}{2}\left(1\pm\sqrt{1-4c^{6}}\right),\end{split}$ (31)
[…]
The dynamics around $(r_{e},v_{e})$ is less clear. Restricting the vector field (28) over the line $r=r_{e}=0$ implies a dynamic
$\begin{cases}\dot{r}=0\\ \dot{v}=C^{2}-\gamma\alpha v\end{cases}$
Hence, the equilibrium $(r_{e},v_{e})$ has an eigenvalue $a^{e}=-\gamma\alpha$ and a stable manifold coinciding with the axis $r=0$. The other eigenvalue is zero and the associated center-manifold $W_{e}^{c}$ is tangent to the vector $(\alpha,-1)$ (Fig. 1 (C)). Locally,
$W^{c}_{e}\approx\left\{v=\frac{C^{2}}{\gamma\alpha}-\frac{r}{\alpha}\right\},\quad\text{for }r>0\text{ small}.$
All orbits with initial conditions $v(0)<0$ and arbitrarily small $r(0)>0$ must approach $W^{c}_{e}$ before entering the circularization spiral, as shown in panel (C).
Figure 1: Phase space portraits for the regularized Kepler system (eq. 28) for $C=\gamma=1$. (A) The conservative case ($\alpha=0$) showing elliptic (blue), parabolic (green), and hyperbolic (orange) orbits. (B) The system with dissipation ($\alpha=0.5$) showing a captured orbit (blue), the weak-unstable-center manifold (red) from the equilibrium $(r_{e},v_{e})=(0,2)$, and an escaping orbit (orange); the green dot marks the circular equilibrium $(r_{c},v_{c})=(1,0)$. (C) Zoom of (B) near $(r_{e},v_{e})$; stable branches along $r=0$ are shown in blue.
To proceed and understand the dynamics at infinity we perform a Poincaré compactification of the vector field (24) introducing new varibles $(x,y)$ with the coordinate transform
$(r,v)\longmapsto(y,x)\quad\text{where}\quad r=\frac{1}{y},\;\;v=\frac{x}{y},$
in a way that $y=0$ corresponds to $r=\infty$. This corresponds to the $U_{1}$ chart in Dumortier et al. [Dumortier_book], pg. 151.
In the new coordinates, equation (24) becomes
$\begin{cases}\dot{y}=-xy\\ \dot{x}=C^{2}y^{4}-x^{2}-\gamma(1+\alpha x)y^{3}\end{cases}$ (32)
The new system (32) has two equilibria: $(y_{c},x_{c})=(\frac{\gamma}{C^{2}},0)$, corresponding to the circular orbit, and
$r_{\infty}\coloneqq(y_{\infty},x_{\infty})=(0,0).$
As the linear approximation of (32) at $(0,0)$ is null, $(y_{\infty},x_{\infty})$ is a degenerate equilibrium. One may note that all unbounded solutions of the Kepler problem ($\alpha=0$) are homoclinic to $(y_{\infty},x_{\infty})=(0,0)$ (see Fig. 2 (A)).
Figure 2: Phase space portraits for the compactified system (32) for $C=\gamma=1$. (A) The system without dissipation ($\alpha=0$) showing elliptic (blue), parabolic (green), and hyperbolic (orange) orbits. The green dot is the circular orbit equilibrium $(y_{c},x_{c})=(1,0)$. (B) The system with dissipation ($\alpha=0.5$) showing a captured orbit (blue), and an escaping orbit (orange). The curve in red represents the weak-unstable-center manifold from the degenerate equilibrium $(r_{e},v_{e})$ which here is a point at infinity. (C) Zoom of (B) near the degenerate equilibrium $(y_{\infty},x_{\infty})=(0,0)$. The hyperbolic, attracting, and repelling sectors are denoted $S_{h},S_{a}$, and $S_{r}$. The purple dashed lines $\Gamma_{u(s)}$ are local approximations to the unstable (stable) manifolds of $(y_{\infty},x_{\infty})=(0,0)$.
To analyze the infinity we blow-up the vector-field at $(y,v)=(0,0)$ using [Dumortier_book, Sec. 3.3]
$(y,x)\longmapsto(p,q)\quad\text{where}\quad y=q^{2},\;\;x=pq^{3}.$
In the new coordinates the vector field (32), after a multiplication by $2q^{-3}$, becomes
$\begin{cases}\dot{q}=-qp\\ \dot{p}=-2\gamma+2C^{2}q^{2}-2\alpha\gamma pq^{3}+p^{2}\end{cases}$ (33)
The $p$-axis, which corresponds to $q^{2}=y=\frac{1}{r}=0$, is invariant. There are two hyperbolic singularities on this axis:
$(q_{\infty+},p_{\infty+})=\left(0,\sqrt{2\gamma}\right),\quad\text{and}\quad(q_{\infty-},p_{\infty-})=\left(0,-\sqrt{2\gamma}\right),$
connected by a solution from $(q_{\infty+},p_{\infty+})$ to $(q_{\infty-},p_{\infty-})$. The whole axis $p$ comes from the blow-up of the equilibrium $(y_{\infty},x_{\infty})$.
The equilibrium $(q_{\infty+},p_{\infty+})$ has a stable manifold $\Gamma_{s}$ that, up to linear approximation, is perpendicular to the $p$-axis: $\Gamma_{s}\approx\{q,p=\sqrt{2\gamma}\}$ for $|q|$ small. Similarly, $(q_{\infty-},p_{\infty-})$ has an unstable manifold $\Gamma_{u}$ that, up to linear approximation, is perpendicular to the $p$-axis: $\Gamma_{u}\approx\{q,p=-\sqrt{2\gamma}\}$ for $|q|$ small.
In the original variables ($y,x$), we then have a local expression for the manifold branches
$\Gamma_{s}=\sqrt{2\gamma}\,y^{\frac{3}{2}}\,,\qquad\Gamma_{u}=-\sqrt{2\gamma}\,y^{\frac{3}{2}}\,,\qquad y>0,$ (34)
which does not depend on $C$ or $\alpha$.
For $\alpha=0$, the vector field (32) has the first integral $\mathcal{H}=\frac{y^{3}\left(C^{2}y-2\gamma\right)+x^{2}}{2y^{2}}$ (from eq. 25), and both characteristic curves $\Gamma_{u}$ and $\Gamma_{s}$ are determined by the equation $\mathcal{H}=0$. One can check that this leads to the approximation in equation (34).
In Fig. 2 we show the local structure of the orbits of the vector field (32) for some values of $\alpha$. The stable curve $\Gamma_{s}$ approaching $(y_{\infty},x_{\infty})$ for positive time, and the curve $\Gamma_{u}$ approaching it for negative time divide the phase space xxyy into three sectors: a hyperbolic sector $S_{h}$, an attracting sector $S_{a}$, and a repelling sector $S_{r}$ (see Fig. 2 (C)).
In the Kepler problem, orbits with positive energy are asymptotically free, with $\mathcal{H}\to\frac{v^{2}}{2}$ as $t\to\pm\infty$. We define scattering orbits, in the dissipative or conservative scenarios, as those for which the following two limits exist
$\lim_{t\to-\infty}v(t)=-v_{\psi}<0\quad\text{and}\quad\lim_{t\to\infty}v(t)=v_{\omega}>0.$
In order to compute the asymptotic shape of a scattering orbit close to a point at infinity, one can use that
$\lim_{t\to\pm\infty}\mathcal{D}=\lim_{t\to\pm\infty}\frac{\alpha}{2}\frac{\gamma}{r^{3}}v^{2}=0$
to neglect dissipation of energy and solve
$\mathcal{H}=\frac{y^{3}\left(C^{2}y-2\gamma\right)+x^{2}}{2y^{2}}=\frac{v^{2}}{2},\quad\text{or}\quad y^{3}\left(C^{2}y-2\gamma\right)+x^{2}=v^{2}y^{2},$
near $(y,x)=(y_{\infty},x_{\infty})=(0,0)$. Up to leading order, the solution is
$x=\pm vy+\mathcal{O}(y^{3/2}).$
Therefore, a scattering orbit leaves the point at infinity tangent to the line $x=-v_{\psi}y$ through the repelling sector $S_{r}$, and arrives at the same point tangent to the line $x=v_{\omega}y$ through the attracting sector $S_{a}$ (see Fig. 2 (C)). In the presence of dissipation, $|v_{\omega}|<|v_{\psi}|$.
To understand the dynamics at the remaning part of infinity, we will make use of a second chart to cover the points $(r,v)=(0,\pm\infty)$ (the $U_{2}$ chart in Dumortier [Dumortier_book], pg. 152). The change of variables from (24) in this case is
$(r,v)\mapsto(w,z)\quad\text{where}\quad w=\frac{r}{v},\quad z=\frac{1}{v}$ (35)
In these coordinates, $z=0$ corresponds to $v=\pm\infty$. At first we will be interested on the limit as $v\to\infty$ that corresponds to the quadrant $w>0$, $z>0$.
After multiplying by $w^{3}$, the vector field (24) in the new chart becomes
$\begin{cases}\dot{w}=w\left(w^{2}+z^{3}\left(\gamma\left(w+\alpha\right)-C^{2}z\right)\right)\\ \dot{z}=z^{4}\left(\gamma\left(w+\alpha\right)-C^{2}z\right)\end{cases}$ (36)
which has two equilibria: $(w_{e},z_{e})=(0,\frac{\alpha\gamma}{C^{2}})$, which corresponds to the $(r_{e},v_{e})$ equilibrium previously found, and a new equilibrium
$v_{\infty}:=(w_{v\infty},z_{v\infty})=(0,0).$ (37)
The circular equilibrium $(r_{c},v_{c})$ is mapped to infinity in this chart.
In Fig. 3 we show phase portraits for the system with and without dissipation. The correctness of this numerically computed phase portrait is corroborated by the following analysis.
We blow up the vector field at $(w_{v\infty},z_{v\infty})=(0,0)$ using [Dumortier_book, Sec. 3.3]
$(w,z)\longmapsto(p,q)\quad\text{where}\quad w=pq^{3},\;\;z=q^{2},$ (38)
to obtain, after division by $q^{6}/2$,
[…]
In the region $p\geq 0$, $q\geq 0$, this equation has three equilibria: $(p,q)=(0,0)$, $(p_{e},q_{e})=(0,\sqrt{\frac{\alpha\gamma}{C^{2}}})$, which corresponds to the ejection point, and a new one $(p,q)=(\sqrt{\frac{\alpha\gamma}{2}},0)$.
Both equilibria $(p,q)=(0,0)$ and $(p,q)=\left(\sqrt{\frac{\alpha\gamma}{2}},0\right)$ are hyperbolic. The equilibrium $(p,q)=(0,0)$ is a saddle, with its unstable manifold lying on the $q$-axis and its stable manifold along the $p$-axis. The equilibrium $(p,q)=\left(\sqrt{\frac{\alpha\gamma}{2}},0\right)$ is a source with strong unstable manifold on the $p$-axis (eigenvalue $2\alpha\gamma$) and weak unstable manifold locally given by $W^{ws}_{\infty}\approx\{p=\sqrt{\frac{\alpha\gamma}{2}},q\}$ for $|q|$ small (eigenvalue $\alpha\gamma$).
In the variables $(w,z)$, the positive branch of $W^{ws}_{\infty}$ is given by
$W^{ws}_{\infty}\approx\left\{w=\sqrt{\frac{\alpha\gamma}{2}}z^{3/2}\right\}\,,\qquad z>0\ \ \text{small},$
while in the original $(r,v)$ variables it corresponds to
$W^{ws}_{\infty}\approx\left\{v=\frac{\alpha\gamma}{2}\frac{1}{r^{2}}\right\}\,,\qquad r>0\ \ \text{small}.$
Figure 3: Phase space portraits for the compactified system (36) for $C=\gamma=1$. (A) The system without dissipation ($\alpha=0$). Dashed lines show the stable and unstable manifolds for $(w_{v\infty},z_{v\infty})=(0,0)$ (blue point) and solid lines a few illustrative orbits. (B) The system with dissipation ($\alpha=0.5$). The equilibrium $(w_{e},z_{e})=(0,0.5)$ (red point) divides the phase space, as indicated by its manifolds (dotted pink lines).
Finally, we analyze the limit $v\to$-$\infty$ that, after the change of variables in equation (35), corresponds to the quadrant $w\leq 0$, $z\leq 0$ in equation (36). We denote the equilibrium at $v=$-$\infty$ as
$(w_{v-\infty},z_{v-\infty})=(0,0)$
A desingularization of $(w_{v-\infty},z_{v-\infty})$ is achieved using variables similar to those in equation (38)
$(w,z)\longmapsto(p,q)\quad\text{where}\quad w=pq^{3},\;\;z=-q^{2}$ (40)
We are interested in the quadrant $p\leq 0$, $q\geq 0$.
In the new variables, the vector field in equation (36) becomes, after division by $q^{6}/2$,
[…]
This equation has a single equilibrium $(p,q)=(0,0)$, which is a saddle. Its unstable manifold lies on the $p$-axis, and its stable manifold lies on the $q$-axis. As expected, the outcome of this analysis is the phase space portrait shown in the quadrant $w\leq 0$, $z\leq 0$ in Fig. 3  [^2] An information provided by the blow up is that in the quadrant $w\leq 0$, $z\leq 0$ solutions initially close to the equilibrium $(0,0)$ approach it approximately along straight lines $z(t)\approx\frac{z(0)}{w(0)}w(t)$ before diverging along the $w$-axis. This can be observed in Fig. 3..
Figure 4: Sketches of the flow line topology on the Poincaré disk for the Kepler system. Green dots mark the circular equilibrium $(r_{c},v_{c})$. (A) Conservative case ($\alpha=0$ or $c=\infty$), showing the parabolic separatrix (pink), which corresponds to the merging of the stable $\Gamma_{s}$ and unstable $\Gamma_{u}$ curves of the “equilibrium” $r_{\infty}$, see equation (4.1). (B) Weakly dissipative case ($\alpha=0.1$ or $c\approx 2.15$), showing that $\Gamma_{s}$ (pink) is homoclinic to $r_{\infty}$, while $\Gamma_{u}$ (green) is attracted to $(r_{c},v_{c})$. All points inside $\Gamma_{s}$ converge to the circular orbit as $t\to\infty$ and diverge to $r=\infty$ as $t\to$-$\infty$. The center-unstable manifold $W^{c}_{e}$ (red) of the ejection point $v_{e}$ converges to $r_{\infty}$ as $t\to\infty$. Orbits outside $\Gamma_{s}$ and below $W^{c}_{e}$ belong to the scattering region, being bi-asymptotic to $r_{\infty}$ as $t\to\pm\infty$. Orbits above $W^{c}_{e}$ form the escaping region, being asymptotic to $r_{\infty}$ as $t\to\infty$ and to $v_{\infty}$, as $t\to$-$\infty$, see equation (37). (C) Strongly dissipative case ($\alpha=0.5$ or $c\approx 1.26$), showing that $\Gamma_{s}$ (pink) is asymptotic to $v_{\infty}$ as $t\to$-$\infty$. Orbits above $\Gamma_{s}$ belong to the escaping region, while those below $\Gamma_{s}$ lie in the basin of attraction of the circular orbit. The center-unstable manifold $W^{c}_{e}$ (red) is attracted to the circular orbit as well as $\Gamma_{u}$ (green). Orbits below $\Gamma_{s}$ and “above” $W^{c}_{e}\cup\Gamma_{u}$ are asymptotic to $v_{\infty}$ as $t\to$-$\infty$. Orbits “below” $W^{c}_{e}\cup\Gamma_{u}$ are asymptotic to $r_{\infty}$ as $t\to$-$\infty$. There exists a critical value $\alpha_{c}\in(0.1,0.5)$ at which $W^{c}_{e}$ merges with $\Gamma_{s}$; at this bifurcation, the set of scattering orbits (present for $0\leq\alpha<\alpha_{c}$) disappears.
4.1.1 Phase portrait topology
To summarize, we recall that the initial vector field (24) describing the radial dynamics of the Kepler problem was regularized via a time rescaling (multiplication by $r^{3}$) yielding vector field (28). Since this vector field is polynomial, it admits a Poincaré compactification onto the disk, where only the half-disk with $r>0$ is relevant to the analysis of the dissipative Kepler problem. The phase portrait topology of this half-disk was assembled from three different charts, each one obtained via local projections by compactification transforms.
To build the full phase portrait topology on the Poincaré disk, we considered three scenarios: $\alpha=0$, $\alpha=0.1$, and $\alpha=0.5$, all for $C=\gamma=1$ (see Fig. 4). In terms of the non-dimensional parameter of the system $c=\frac{C}{\gamma^{2/3}\alpha^{1/3}}$, equation (27), these parameters correspond respectively to: $c=\infty$, $c\approx 2.15$, and $c\approx 1.26$. We remark that the basin of attraction of $r_{c}$ increases along with $\alpha$, as expected.
4.2 The angular-momentum-energy space
The angular-momentum-energy map, or CH map, is defined as the map that associates to each state $(\textbf{r},\textbf{v})\in\mathbb{R}^{2}/\{0\}\times\mathbb{R}^{2}$ of the dissipative Kepler problem the pair $(C=(\textbf{r}\times\textbf{v})\cdot\hat{\textbf{z}},\mathcal{H})\in\mathbb{R}^{2}$. For a given $C$, the minimum of $\mathcal{H}=\frac{v^{2}}{2}+\frac{C^{2}}{2r^{2}}-\frac{\gamma}{r}$ is achieved at the circular orbit, $r_{c}=\frac{C^{2}}{\gamma}$ and $v=0$, with energy $\mathcal{H}_{\text{min}}=-\frac{\gamma^{2}}{2C^{2}}$. Thus, the image of the CH map is the region above the graph in Fig. 5.
Figure 5: Curve of minimum energy $\mathcal{H}$ as a function of the angular momentum $C$, achieved at the circular orbit. Every orbit is mapped as a segment on the vertical lines above this curve. In green, an orbit captured by the circular attractor. In red, an example of a scattered (escape) orbit.
An orbit in the conservative Kepler system maps to a single point in the CH space, as $C$ and $\mathcal{H}$ are constants. In the presence of dissipation, however, the image of an initial point, not on the curve of minimum energy, always moves vertically down. There are two possibilities for this motion as $t\to\infty$. One, where the point moves vertically down till it reaches the minimum energy curve, which happens if, and only if, the initial point is in the basis of attraction of the circular orbit. The second possibility is that the point moves vertically down until reaching a limit with $\mathcal{H}>0$. This scenario happens if, and only if, the initial point belongs either to the scattering region or to the escaping region.
4.3 Collisions and almost collisions.
In the dissipative Kepler problem considered here, collisions differ significantly from those in the conservative case.
Collisions occur only when the angular momentum CC is zero. In this case, independently of the value of the dissipative parameter $\alpha\geq 0$, the orbits lie on straight segments with the origin, which is the collision point, as an endpoint.
In the conservative case, an almost collision corresponds to an elliptic orbit with eccentricity $e$ close to one and one focus at the origin. Since $C=\sqrt{\gamma a(1-e^{2})}$, where $a$ is the semimajor axis, we have $C\approx 0$. The orbit then approximates a segment of length $2a$, with one endpoint (the periapsis) close to the origin. This suggests a collision regularization procedure: for $C=0$, the orbit becomes a segment of length $2a$ with the origin as one endpoint, and the particle moves periodically between the endpoints with infinite velocity at the periapsis.
In the following paragraphs we show that the dynamical behavior of collision orbits changes qualitatively when $\alpha>0$, regardless of how small $\alpha$ is.
This effect is a consequence of the specific structure of the dissipative force. The three parameters in equation (24) can be reduced to the single parameter $c=\frac{C}{\gamma^{2/3}\alpha^{1/3}}$, given in equation (27), by choosing $\alpha$ as the unit of time and $\gamma^{1/3}\alpha^{2/3}$ as the unit of length. With this scaling, system (24) becomes
$\begin{cases}\dot{r}=v\\ \dot{v}=\frac{c^{2}}{r^{3}}-\left(\frac{1}{r^{2}}+\frac{v}{r^{3}}\right)\end{cases}$ (42)
If $\alpha\ll 1$, the time $\tau$ in equation (42) is a slow time, related to the original time $t$ by $t=\alpha\tau$, while the length scale $\lambda$ corresponds to a magnified spatial variable $\ell=\gamma^{1/3}\alpha^{2/3}\lambda$. Velocities are correspondingly rescaled as $\frac{\Delta\ell}{\Delta t}=\frac{\gamma^{1/3}}{\alpha^{1/3}}\frac{\Delta\lambda}{\Delta\tau}$.
The key point is that, for $\alpha>0$, the limit $C\to 0$ (equivalently $c\to 0$) in equation (42) produces the same effective dynamics as the strong dissipative limit obtained by keeping $C>0$ fixed and letting $\alpha\to\infty$. Thus, in this model, close approaches between the bodies are intrinsically associated with strong energy dissipation.
We now return to the original time and length scales used in equation (24). For $\alpha>0$ and $C>0$ sufficiently small so that $c=\frac{C}{\gamma^{2/3}\alpha^{1/3}}$ is small, most solutions are attracted to the circular periodic orbit, as shown in Figure 4. In this regime, the radius $r_{c}=\frac{C^{2}}{\gamma}$ is small, and the motion rapidly settles into a circular orbit close to the origin.
For $c^{6}<\frac{1}{4}$, or $c\approx 0.79$, see equation (31), $(r_{c},0)$ is an attracting node. Thus, $r(t)$ with $r(0)>r_{c}$ converges to $r_{c}$ with at most one local maximum when $\dot{r}(0)>0$.
In contrast, for $\alpha=0$ and fixed small $C>0$, the motion is elliptical with eccentricity close to one. If $\alpha>0$ is introduced while keeping $c=\frac{C}{\gamma^{2/3}\alpha^{1/3}}\gg 1$, the ellipse slowly shrinks toward the circular orbit of radius $r_{c}=\frac{C^{2}}{\gamma}$. Increasing $\alpha$ further so that $c<0.79$, or $\alpha>\frac{2C^{3}}{\gamma^{2}}$, leads to rapid convergence: most of the orbital energy is dissipated near the first perihelion passage, and $r(t)\to r_{c}$ with at most one oscillation. Note that as $C\to 0$, the threshold $\alpha>\frac{2C^{3}}{\gamma^{2}}$ also tends to zero.
At the same time, as $C\to 0$, the ejection equilibrium $(0,v_{e})=\left(0,\frac{C^{2}}{\gamma\alpha}\right)$ given in equation (30) converges to $(0,0)$. Hence, as $C\to 0$, $(0,v_{e})$ approaches $(r_{c},0)$, and these equilibria merge at $C=0$.
We now analyze the collision orbits, $C=0$, for any fixed value of $\alpha>0$. In this case, equation (42) becomes
$\begin{cases}\dot{r}=v\\ \dot{v}=-\left(\frac{1}{r^{2}}+\frac{v}{r^{3}}\right)\end{cases}$ (43)
As in equation (28), we regularize the vector field in equation (43) by multiplying it by $r^{3}$:
$\begin{cases}\dot{r}=r^{3}v\\ \dot{v}=-\left(r+v\right).\end{cases}$ (44)
Equation (44) has a single equilibrium $(r,v)=(0,0)$, with eigenvalues 0 and 1-1. This equilibrium admits an attracting center manifold tangent to the line v=rv=-r and a stable manifold tangent to the $v$-axis. A computation shows that the center manifold is given by
$v=h(r)=-r-r^{4}(1+5r^{3}+44r^{6}+539r^{9}+8337r^{12}+154632r^{15}\ldots)$ (45)
Since $h(r)<0$ for $r>0$, equation (44) implies that on the center manifold $\dot{r}=r^{3}h(r)<0$, so $(r,v)=(0,0)$, which represents collisions, is attracting.
The transverse attraction to the center manifold is stronger than the attraction along it; thus, solutions rapidly approach the center manifold and then drift along it toward $r=0$. Figure 6 shows the approximation of the center manifold (green) together with representative trajectories of (44).
Since the graph $v=h(r)$ is also invariant under the flow of equation (43), in the dissipative case $\alpha>0$ (with $\alpha$ normalized to one and length scale $\gamma^{1/3}\alpha^{2/3}$ equal to one), collisions occur with finite speed approximately given by equation (45), as $\dot{r}=v$.
Figure 6: The center manifold of $(r,v)=(0,0)$ as given in equation (45) and several stream lines of the vector field (44). The topology of the stream lines of equation (43) is the same.
In conclusion, while in the conservative case regularized collision orbits oscillate along a segment with infinite velocity at the origin, in the dissipative case, for any $\alpha>0$, a collision orbit with $\dot{r}(0)<0$ moves monotonically toward the origin with finite velocity.
5 Averaging
The goal in this Section is to average equations (24) with respect to the mean anomaly of the Kepler problem. We start by writing equations (24) in an intrinsic way as
$i_{X}\omega_{s}=dH+Q$ (46)
where $X$ is the dissipative vector field. In the canonical set of variables $(r,\theta,\dot{r},C)$ used in Section 4, the symplectic form is
$\omega_{s}=dr\wedge d\dot{r}+d\theta\wedge dC,$
and $Q$, which we will call dissipative one-form, is given by
$Q=\alpha\gamma\frac{\dot{r}}{r^{\beta}}dr=\frac{\alpha\gamma}{2}\frac{\dot{r}}{r^{\beta+1}}d(r^{2}).$ (47)
The angle $\theta$ is the true longitude that can be decomposed as
$\theta=\varpi+f$
with $f$ as the true anomaly and $\varpi$ as the longitude of the periapsis.
There is a canonical change of variables from $(r,\theta,\dot{r},C)$ to the Delaunay variables [Laskar2017] given by
[…]
where: $a$ is the semimajor axis, $M$ is the mean anomaly, and $\Lambda$ is the Delaunay action conjugated to $M$. In Delaunay’s variables
$\omega_{s}=dM\wedge d\Lambda+d\varpi\wedge dC\,,\qquad H=-\frac{\gamma^{2}}{2\Lambda^{2}}$
It remains to write the dissipative form $Q=\alpha\gamma\frac{\dot{r}}{r^{3}}dr$ in terms of the Delaunay variables.
In order to write the variables $r$ and $\dot{r}$ in terms of $M$, $\Lambda$, and the eccentricity $e$ we will use the Hansen’s expansions given by:
$\left(\frac{r}{a}\right)^{n}{\rm e}^{imf}=\sum_{k=-\infty}^{\infty}X^{n,m}_{k}(e){\rm e}^{ikM}\,$, (49)
where $n$, $m$, and $k$ are integers and $e$ is the eccentricity.  [^3] The Hansen coefficients have the following properties: $X^{n,m}_{k}\in\mathbb{R}\,,\qquad X^{n,m}_{-k}=X^{n,-m}_{k}$ (50) and $X^{n,m}_{k}(e)=e^{|m-k|}G^{n,m}_{k}(e^{2})$ (51) where $G^{n,m}_{k}$ is analytic for $e^{2}<1$.
The eccentricity can be written in terms of $\Lambda$ and $C$, and therefore $Q$ can be expressed in the Delaunay variables. However, the expression for the dissipative form becomes simpler if we replace $C$ in Delaunay’s variables by the eccentricity $e$. The set of variables $(M,\varpi,\Lambda,e)$ is not canonical, so we must compute the symplectic form in this coordinates.
From equation (48) we obtain
$dC=\sqrt{1-e^{2}}d\Lambda-\frac{e\Lambda}{\sqrt{1-e^{2}}}de=\eta d\Lambda-\frac{e\Lambda}{\eta}de$
In the variables $(M,\varpi,\Lambda,e)$ the symplectic form becomes
$\omega_{s}=dM\wedge d\Lambda+\eta\,d\varpi\wedge d\Lambda-\frac{e\Lambda}{\eta}d\varpi\wedge de$
As a result the dissipative Hamiltonian equations become:
$\left(\begin{array}[]{l}\dot{M}\\ \dot{\varpi}\\ \dot{\Lambda}\\ \dot{e}\end{array}\right)=\underbrace{\left(\begin{array}[]{cccc}0&0&1&\frac{\eta^{2}}{\Lambda e}\\ 0&0&0&-\frac{\eta}{\Lambda e}\\ -1&0&0&0\\ -\frac{\eta^{2}}{\Lambda e}&\frac{\eta}{\Lambda e}&0&0\end{array}\right)}_{\textbf{S}}\left(\begin{array}[]{c}\partial_{M}H+Q_{M}\\ \partial_{\varpi}H+Q_{\varpi}\\ \partial_{\Lambda}H+Q_{\Lambda}\\ \partial_{e}H+Q_{e}\end{array}\right)$ (52)
where $Q=Q_{M}dM+Q_{\varpi}d\varpi+Q_{\Lambda}d\Lambda+Q_{e}de$.
We use the Hansen’s expansion to write [^4] Consider the Kepler vetor field $X$ given on a subset $U\subset\mathcal{R}^{2}\times\mathcal{R}^{2}$, for instance, with coordinates $(\textbf{x},\textbf{v})$, where all orbits are elliptic. The elements $a,e,\varpi,C,r,v,f,\ldots$ are functions defined on $U$. The vector field $X$ acts on these functions as the time derivative, therefore $X\cdot r=\dot{r}=v\,,\quad X\cdot e=X\cdot a=0\,,\quad X\cdot f=\dot{f}\,\ldots$ The Lie derivative of the differential of any function gives: $L_{X}dr=d(X\cdot r)X=d\dot{r}$, so the time derivative commutes with the differential. In the coordinate system $(M,\Lambda)$, the vector field $X$ is given by $X=n\partial_{M}$. We must be cautious in using the notation $X=n\partial_{M}$. For instance, if $F=F(r,M)$ then $\partial_{M}F(r,M)$ can be understood as the partial derivative of $F$ with r=r=constant, which is false, since $\partial_{M}$ means the derivative of the function while $\Lambda$ is constant.
$\frac{\dot{r}}{r^{\beta+1}}=-\frac{d}{dt}\frac{r^{-\beta}}{\beta}=-\frac{1}{\beta}\frac{d}{dt}\left(a^{-\beta}\sum_{j=-\infty}^{\infty}X^{-\beta,0}_{j}{\rm e}^{ijM}\right)$
Since $\dot{a}=\dot{e}=0$, $\dot{M}=n=\sqrt{\gamma}a^{-3/2}$, and $\Lambda=\sqrt{\gamma a}$, we obtain:
$\frac{\dot{r}}{r^{\beta+1}}=-i\frac{\gamma^{\beta+2}}{\beta\Lambda^{2\beta+3}}\sum_{j=-\infty}^{\infty}jX^{-\beta,0}_{j}{\rm e}^{ijM}$ (53)
We again use Hansen’s expansion to write
$\begin{split}dr^{2}=&d\left(a^{2}\sum_{k=-\infty}^{\infty}X^{2,0}_{k}{\rm e}^{ikM}\right)=d\left(\gamma^{-2}\Lambda^{4}\sum_{k=-\infty}^{\infty}X^{2,0}_{k}{\rm e}^{ikM}\right)\\ =&i\gamma^{-2}\Lambda^{4}\sum_{k=-\infty}^{\infty}kX^{2,0}_{k}{\rm e}^{ikM}dM\\ &+4\gamma^{-2}\Lambda^{3}\sum_{k=-\infty}^{\infty}X^{2,0}_{k}{\rm e}^{ikM}d\Lambda\\ &+\gamma^{-2}\Lambda^{4}\sum_{k=-\infty}^{\infty}\big(X^{2,0}_{k}\big)^{\prime}{\rm e}^{ikM}de\end{split}$ (54)
where $\big(X^{2,0}_{k}\big)^{\prime}=\frac{d}{de}\big(X^{2,0}_{k}\big)$.
Therefore, from equations (47), (53), and (54) we obtain
$\begin{split}dQ=&\frac{\alpha\gamma^{\beta+1}}{2\beta\Lambda^{2\beta-1}}\sum_{j=-\infty}^{\infty}\sum_{k=-\infty}^{\infty}jkX^{-\beta,0}_{j}X^{2,0}_{k}{\rm e}^{i(k+j)M}dM\\ &-2i\frac{\alpha\gamma^{\beta+1}}{\beta\Lambda^{2\beta}}\sum_{j=-\infty}^{\infty}\sum_{k=-\infty}^{\infty}jX^{-\beta,0}_{j}X^{2,0}_{k}{\rm e}^{i(k+j)M}d\Lambda\\ &-i\frac{\alpha\gamma^{\beta+1}}{2\beta\Lambda^{2\beta-1}}\sum_{j=-\infty}^{\infty}\sum_{k=-\infty}^{\infty}jX^{-\beta,0}_{j}\big(X^{2,0}_{k}\big)^{\prime}{\rm e}^{i(k+j)M}de\end{split}$ (55)
For $\alpha=0$, equations (52) become $\dot{M}=n>0$ and $\dot{\varpi}=\dot{\Lambda}=\dot{e}=0$. This allows us to average the equations with respect to $M$.
We define the averaging operator
$\langle h\rangle=\frac{1}{2\pi}\int_{0}^{2\pi}h(M)dM$ (56)
Since the matrix S in equation (52) does not depend on the mean anomaly the averaged equations become
$\left(\begin{array}[]{l}\dot{M}\\ \dot{\varpi}\\ \dot{\Lambda}\\ \dot{e}\end{array}\right)=\left(\begin{array}[]{cccc}0&0&1&\frac{\eta^{2}}{\Lambda e}\\ 0&0&0&-\frac{\eta}{\Lambda e}\\ -1&0&0&0\\ -\frac{\eta^{2}}{\Lambda e}&\frac{\eta}{\Lambda e}&0&0\end{array}\right)\left(\begin{array}[]{c}\langle Q_{M}\rangle\\ \langle Q_{\varpi}\rangle\\ \partial_{\Lambda}H+\langle Q_{\Lambda}\rangle\\ \langle Q_{e}\rangle\end{array}\right)$ (57)
Equation (55) and $X^{n,0}_{-k}=X^{n,0}_{k}$ imply that the averaged components of the dissipative form are:
$\begin{split}\langle Q_{M}\rangle=&-\frac{\alpha\gamma^{\beta+1}}{2\beta\Lambda^{2\beta-1}}\sum_{k=-\infty}^{\infty}k^{2}X^{-\beta,0}_{k}X^{2,0}_{k}\\ \langle Q_{\varpi}\rangle=&0\\ \langle Q_{\Lambda}\rangle=&2i\frac{\alpha\gamma^{\beta+1}}{\beta\Lambda^{2\beta}}\underbrace{\sum_{k=-\infty}^{\infty}kX^{-\beta,0}_{k}X^{2,0}_{k}}_{=0}\\ \langle Q_{e}\rangle=&i\frac{\alpha\gamma^{\beta+1}}{2\beta\Lambda^{2\beta-1}}\underbrace{\sum_{k=-\infty}^{\infty}kX^{-\beta,0}_{k}\big(X^{2,0}_{k}\big)^{\prime}}_{=0}\end{split}$ (58)
Remark
$\langle Q_{M}\rangle=\frac{1}{n}\underbrace{\langle 2\mathcal{D}\rangle}_{\langle\dot{E}\rangle}=\Delta E\geq 0$ (59)
where
$\Delta E(a,e)=\frac{T}{2\pi}\int_{0}^{2\pi}\dot{E}dM=\int_{0}^{T}\dot{E}dt$ (60)
is the energy dissipated in one period along an elliptic orbit with semi-axis $a$ and eccentricity $e$.
Indeed
$\begin{split}\frac{(\textbf{v}\cdot\hat{\textbf{r}})^{2}}{|\textbf{r}|^{\beta}}&=\frac{\dot{r}^{2}}{r^{\beta}}=-\frac{1}{2\beta}\frac{dr^{2}}{dt}\frac{dr^{-\beta}}{dt}=-\frac{1}{2\beta a^{\beta-2}}\frac{dr^{2}a^{-2}}{dt}\frac{dr^{-\beta}a^{\beta}}{dt}\\ &=-\frac{1}{2\beta a^{\beta-2}}\frac{d}{dt}\left(\sum_{j}X^{2,0}_{j}{\rm e}^{ijM}\right)\frac{d}{dt}\left(\sum_{k}X^{-\beta,0}_{k}{\rm e}^{ikM}\right)\\ &=\frac{n^{2}}{2\beta a^{\beta-2}}\sum_{j\,,k}X^{2,0}_{j}X^{-\beta,0}_{k}jk{\rm e}^{i(j+k)M}\end{split}$
Thus
$\frac{1}{2\pi}\int_{0}^{2\pi}\frac{(\textbf{v}\cdot\hat{\textbf{r}})^{2}}{|\textbf{r}|^{\beta}}dM=-\frac{n^{2}}{2\beta a^{\beta-2}}\sum_{k}X^{2,0}_{k}X^{-\beta,0}_{k}k^{2}$
The result follows from equations (47) and (58).
In Figure 7 we show three plots of the function $\sum_{k}k^{2}X^{2,0}_{k}(e)X^{-\beta,0}_{k}(e)$ for $\beta=2,3,8$. For constant values of $a$ and $e$, this figure shows that the dissipation is stronger when $\beta$ increases.
Figure 7: The function $F_{\beta}(e)=-\sum_{k}k^{2}X^{2,0}_{k}(e)X^{-\beta,0}_{k}(e)$ for $\beta=2,3,8$. The following expansions hold: $F_{2}(e)=e^{2}+O\left(e^{4}\right)$, $F_{3}(e)=\frac{3e^{2}}{2}+O\left(e^{4}\right)$, and $F_{8}(e)=4e^{2}+O\left(e^{4}\right)$.
The averaged equations are $\dot{M}=n$, $\dot{\varpi}=0$ and
$\begin{split}\dot{\Lambda}&=-\langle Q_{M}\rangle<0\\ \dot{e}&=-\frac{\eta^{2}}{\Lambda e}\langle Q_{M}\rangle<0\end{split}$ (61)
Using $C=\Lambda\eta$ we obtain that $C$ remains constant.
Note that $e=0$ is an attracting equilibrium of the equation for $\dot{e}$. As $e\to 0$, the constancy of the angular momentum $C=\Lambda\sqrt{1-e^{2}}$ implies that $\Lambda\to C$. Therefore $a\to C^{2}/\gamma$ and the asymptotic value of $a$ is determined by the initial angular momentum.
The Gauss form of the third Kepler law, namely,
$a^{3}n^{2}=\frac{C^{2}}{\tilde{p}}=m$
where $\tilde{p}=a(1-e^{2})=a\eta^{2}$ is the semi-latus rectum, implies that the semi-latus rectum remains constant under the action of the dissipative force.
One may note that the semi-major axis is always decreasing, so the model cannot explain the Earth-Moon increasing distance but it can explain the Mercury-Sun decreasing distance.
6 Conclusion
We sought to understand the effect of a dissipative force that while removing energy preserves angular momentum. The force was motivated by a simplified tidal model for punctual masses, where the velocity dependence is linear and projected radially, making the force central. This form for the force preserves rotational symmetry while breaking time symmetry, and is trivially extendable to systems with any number of bodies.
In particular, we showed that a radial force whose dependence is homogeneous of degree $d=3$ yields an extra ‘radial’ symmetry matching that of gravity. This allows us to easily derive the equations of motion for planar central configurations for generic $N$ bodies, with the resulting system for its homographic parameters being equivalent to the dissipative Kepler problem.
The global phase space portrait for the dissipative Kepler problem was obtained via a Poincaré compactification of the regularized vector field. The technique used relied on building three different charts to cover the Poincaré disk, where the dynamics near infinity can be analyzed locally.
The topology of flow lines in the dissipative Kepler system undergoes a transition as the dissipation strength increases, where the separatrix of bounded (captured) and unbounded (scattering) orbits in the conservative case reconnects with the weak-center-manifold (from $(r_{e},v_{e})$) linking to escaping orbits. This reconnection creates a new division between captured and scattered orbits, where all approaches between bodies are bounded and lead to circularization.
Collision orbits in the conservative and dissipative settings are fundamentally different. While in the conservative case regularized collision orbits oscillate along a segment with infinite velocity at the origin, in the dissipative case, for any $\alpha>0$, a collision orbit with $\dot{r}(0)<0$ moves monotonically toward the origin with finite velocity.
As mentioned above, the Kepler system corresponds to the homographic dynamics of any planar central configuration under dissipation. For sufficiently small dissipation and nonzero angular momentum, configurations in the basins of attraction of the circular equilibrium have their homothetic (radial) motion damped, leading to purely rotational motion and hence to a relative equilibrium. In the dissipative case, purely homothetic solutions lead to collisions with finite velocity, as described above and in more detail in Section 4.3.
To extend the current model, one could consider the effect of rotation of bodies, reintroducing some of the neglected tidal effects. Moreover, to increase flexibility, one may also assign different dissipation strengths amongst pairs of bodies, $\alpha\to\alpha_{ij}=\alpha_{i}\alpha_{j}$, adapting to realistic scenarios where the primary bodies respond differently to tides due to its inner structure.
[…]
Acknowledgements
This research was funded by the São Paulo Research Foundation (FAPESP) under grants 2024/15052-0 (M.L.) and 2023/07076-4 (C. R.).
C.R. thanks several colleagues with whom he discussed forces similar to the one considered in this paper: Claudio Asano, Lucas Ruiz, and especially Antonio Fernandes, who performed several computations on the dissipative Kepler problem in a different way that helped in understanding the problem.
Finally, we would like to acknowledge Prof. Mark Levi, who informed us about the PhD thesis of Chong Ai and for many interesting conversations on the subject of this paper and related topics.